OpenStudy (anonymous):

equation y = 2 x^2 as a polar equation

zepdrix (zepdrix):

Use these when converting from Cartesian to Polar:\[\Large y=r \sin \theta \qquad\qquad\qquad x=r \cos \theta\]Plugging in gives us,\[\Large y=2(x)^2 \qquad\to\qquad r \sin \theta=2(r \cos \theta)^2\]

zepdrix (zepdrix):

Do you know how to solve for \(\large r\) from that point? :D

OpenStudy (anonymous):

yes! thanks so much!!

OpenStudy (loser66):

@zepdrix he knows but I don't, please show me.

zepdrix (zepdrix):

Show us maria!! XD

OpenStudy (anonymous):

rsin(theta)=2x^2 rsin(theta)=rcos^2(theta) tan(theta)=2rcos(theta) r=(sec(theta)tan(theta))/2 :)